OpenStudy (anonymous):

There are 10 people at a meeting, some men and some women. The total number of men is 2 more than the number of women. How many men and women are at the party? In order to graph, let the number of men be x and the number of women be y.

OpenStudy (anonymous):

4women 6men

OpenStudy (anonymous):

I need this in slope intercept form so that I may graph it.

OpenStudy (anonymous):

Say x is women, y is men x+y=10 y=x+2 plug in the the second equation into the first to get x+(x+2)=10 then solve for x. Use this value of x to solve for y in the second equation

OpenStudy (anonymous):

I'm sorry to be so slow at this but I'm still not understanding what to do with these equations.

OpenStudy (anonymous):

x is number of man, y number of woman, number of man is 2> number of woman that's mean x=y+2, then you have x+y=10=>( y+2)+y=10=> 2y+2=10 or 2y=8 so y=4, x=y+2 then x=6

OpenStudy (anonymous):

for the first you have to make 2 equations the first x+y=10 and x=y+2 than it's all simple :)

OpenStudy (anonymous):

Men = x. Women = y. The total number of people (10) = women (y) + men (x) 10 = y + x. Because there are two more men than there are women, y (women) = x (men) - 2. So your two equations are y + x = 10 and y = x - 2. Because you know the equation for y (x - 2) you can substitute that in for y in the first equation (y + x = 10). The equation you would end up with is this: (x - 2) + x = 10. 2x - 2 = 10. 2x = 12. x = 6. You can then plug that in for x in the first equation, in order to find y. y + x = 10. y + 6 = 10. y = 4.

OpenStudy (anonymous):

Thank You All So Very Much!

OpenStudy (anonymous):

u r welcome ^_^